OpenStudy (anonymous):

3 x (1/2)^x Can someone go through the domain, range, and horizontal asympote with me? And other problems I have on the same topic

OpenStudy (michele_laino):

is your function this? \[f(x)=3x*\left( \frac{ 1 }{ 2 } \right)^{x}\]

OpenStudy (anonymous):

The 3 is by itself

OpenStudy (michele_laino):

\[f(x)=3*x*\left( \frac{ 1 }{ 2 } \right)^{x}\]

OpenStudy (michele_laino):

is it ok, now?

OpenStudy (anonymous):

\[3\times(1/2)^{x}\]

OpenStudy (michele_laino):

ok! I understand

OpenStudy (michele_laino):

domain of of is the set of real numbers, namely \[\mathbb{R} \]

OpenStudy (michele_laino):

because your function is continuous everywhere in R, do you agree?

OpenStudy (michele_laino):

furthermore, since exponential function is Always positive, also your function is, so range of your function is the set: \[[0,+\infty)\]

OpenStudy (anonymous):

Yes I get it so far

OpenStudy (michele_laino):

Sorry I lost my connection

OpenStudy (michele_laino):

ok! in order to find horizontal asymptotes we have to calculate these limits: \[\lim _{\pm \infty}f(x)=\lim _{\pm \infty }\frac{ 3 }{ 2^{x} }\]

OpenStudy (michele_laino):

now: \[\lim _{x \rightarrow +\infty}f(x)=0\] and: \[\lim _{x \rightarrow -\infty}f(x)=+\infty\] is it ok?

OpenStudy (anonymous):

I don't get that part

OpenStudy (michele_laino):

is it clear for you?

OpenStudy (michele_laino):

for the first limit, keep in mind that when x-->+infinity, then 2^x--->+infinity, so: \[\frac{ 1 }{ 2^{x} }\rightarrow 0\]

OpenStudy (michele_laino):

for second limit, keep in mind that when x-->-infinity, then 2^x-->0 through positive numbers, so: \[\frac{ 1 }{ 2^{x} }\rightarrow + \infty\]

OpenStudy (michele_laino):

now?

OpenStudy (anonymous):

Yes, I think I got it

OpenStudy (michele_laino):

ok! since f(x) is continuous every where, there aren't vertical asymptotes, so we can go to find maximum and minimum points

OpenStudy (michele_laino):

first we note that: \[f(0)=\frac{ 3 }{ 2^{0} }=3\]

OpenStudy (michele_laino):

so we can draw this: 

OpenStudy (michele_laino):

now I calculate f'(x): \[f'(x)=-\frac{ 3 }{ 2^{x} }\ln 2\]

OpenStudy (michele_laino):

now I ask you at which points we have f'(x)=0?

OpenStudy (michele_laino):

or which is the sign of f'(x)?

OpenStudy (anonymous):

I don't really know

OpenStudy (michele_laino):

@minisweet4 since f'(x) is Always negative, in the domain of f(x), we can conclude that f(x) is always decreasing

OpenStudy (michele_laino):

furthermore there aren't any inflection points

OpenStudy (anonymous):

I get it

OpenStudy (michele_laino):

I think that graph of your function is: